diagonalize A= {{2,-1,-1},{1,4,1},{-1,-1,2}}
and compute A^2345
1
Expert's answer
2020-09-03T08:27:12-0400

"A=\begin{pmatrix} 2 & -1 & -1 \\
1 & 4 & 1 \\
-1 & -1 & 2\end{pmatrix} = \begin{pmatrix} 0 & -1 & -1 \\
6 & 4 & 1 \\
0 & -1 & 2\end{pmatrix} = \begin{pmatrix} 0 & -1 & -1 \\
6 & 4 & 1 \\
0 & 0 & 3\end{pmatrix}" =

"\begin{pmatrix} 0 & -1 & -1 \\
6 & 4 & 1 \\
0 & 0 & 3\end{pmatrix} = \begin{pmatrix} 0 & -1 & -1 \\
6 & 3 & 0 \\
0 & 0 & 3\end{pmatrix} = \begin{pmatrix} 6 & 2 & -1 \\
6 & 3 & 0 \\
0 & 0 & 3\end{pmatrix}" =

"= \begin{pmatrix} 6 & 2 & 0 \\
6 & 3 & 0 \\
0 & 0 & 3\end{pmatrix} = \begin{pmatrix} 2 & 2 & 0 \\
0 & 3 & 0 \\
0 & 0 & 3\end{pmatrix} = \begin{pmatrix} 2 & 0 & 0 \\
0& 3 & 0 \\
0 & 0 & 3\end{pmatrix}"

Since A is diagonal now, it is easy to find A to any power.

"A^{2345} =\begin{pmatrix} 2^{2345} & 0 & 0 \\
0& 3^{2345} & 0 \\
0 & 0 & 3^{2345}\end{pmatrix}"